10. Continuous Compounding Compute the future value of $1,625 continuously compounded for a. Five years at an annual percentage rate of 14 percent. b. Three years at an annual percentage rate of 6 percent. c. Ten years at an annual percentage rate of 8 percent d. Eight years at an annual percentage rate of 9 percent CHAPTER

please show all work
Answer #1

Please provide rating...

Formula for contineous compounding future value= P*e^rt
* we have to find the value of e using table of calculator or excel I am using excel .
a b c d
P= 1625 1625 1625 1625
r=interest rate 14% 6% 8% 9%
t=year 5 5 5 5
r*t= 0.70 0.30 0.40 0.45
e^rt 2.013753 1.349859 1.491825 1.568312
Excel formula =exp(0.7) =exp(0.3) =exp(0.4) =exp(0.45)
P*e^rt 3272.35 2193.52 2424.22 2548.51
Therefore Future value = 3272.35 2193.52 2424.22 2548.51